OpenStudy (anonymous):

What is the inverse of f(x) = log8(x)

OpenStudy (campbell_st):

let \[x = \log _{8} y\] and make y the subject \[8^x = y\]

OpenStudy (anonymous):

so inverse = (f-1)x = 8 ^x